# Regret Optimal Control for Uncertain Stochastic Systems

### v2

@@ Abstract
I Introduction
@@
For the case of known cost functions, the formulation of [5, 6, 7] has received increasing interest thanks to: optimality of the clairvoyant benchmark policy, possibility of computing the regret-minimizing controller, and remarkable performance reported in several applications, including longitudinal motion control of a helicopter and control of a wind energy conversion system [8]. In particular, among recent contributions, [9] and [10] proposed an efficient optimization-based synthesis framework to incorporate safety constraints, [11] established recursive feasibility and stability guarantees for receding horizon regret optimal control, [12] and [13] considered partially-observed systems, [14] and [8] investigated the closely related metric of competitive ratio, [7] and [15] considered state estimation problems, and [16] studied connections with imitation learning.
@@
For a wide range of control applications, including robotics, building energy management, and power grids [26], designing a single state feedback policy that attains robust performance across all admissible system dynamics can prove overly conservative. Instead, it is beneficial to optimize for a unique closed-loop behavior – while allowing the state feedback law that achieves it to vary – leveraging a posteriori measurements of exogenous perturbations such as external forces, solar radiation, and electricity demands for control implementation.
@@
II Problem Statement and Preliminaries
II-A Dynamics, control objective, and constraints
@@
Often times, the probability distribution $\mathbb{P}_{\bm{\theta}}$ is unknown, yet uncertainty samples are directly made available to the policy designer as observations. For instance, this is the case when the realizations $\bm{\theta}^{k}\in\mathcal{D}$ correspond to a series of system identification experiments, see, e.g., [18] and [28], or to the values of the physical parameters of a batch of components, which scatter around their nominal value due to tolerance levels and variability in the production process.
@@
As previously discussed, measurable disturbance processes arise in several control applications and include, e.g., reference changes, friction, and external forces in robotics, and heat and humidity loads in heating, ventilating and air conditioning systems. Besides, our formulation encompasses the broad class of linear-parameter-varying systems. In fact, if $\theta_{t}$ denotes an a priori uncertain but measurable scheduling parameter, then past disturbance realizations can always be reconstructed by $w_{t}=E_{t}(\theta_{t})^{\dagger}(x_{t+1}-A_{t}(\theta_{t})x_{t}-B_{t}(\theta_{t})u_{t})$, where $E_{t}(\theta_{t})^{\dagger}$ is the Moore-Penrose inverse of $E_{t}(\theta_{t})$. Differently from most literature on linear-parameter-varying systems, however, we allow generic nonlinear dependence with respect to $\theta_{t}$ of the system matrices $A_{t}(\theta_{t})$, $B_{t}(\theta_{t})$, and $E_{t}(\theta_{t})$.
@@
II-B Linear disturbance feedback policy
In general, it is well-known that optimizing over the function space of feedback policies is computationally intractable. Therefore, as common in the control literature [29, 30], throughout this paper we restrict our attention to linear disturbance feedback policies of the form $\bm{u}=\bm{\Phi}_{u}\bm{w}$, with $\bm{\Phi}_{u}$ lower block-triangular to enforce causality. [^6] By carefully adapting the convex reformulation of Proposition 1 below, our results can be extended to the case of affine policies $\bm{u}=\bm{\Phi}_{u}\bm{w}+\bm{v}$ through the definition of an augmented disturbance vector $\bm{\delta}=(\bm{w},1)$. We note that linear policies enable attaining minimum regret against the optimal sequence of control actions in hindsight if the true system dynamics are known and the safety constraints are not active [5, 6, 7]. Moreover, as we will show in the next section, unlike linear state feedback policies, this choice allows us to approximate the intractable minimization of (4) subject to the dynamics (1) and the constraints (5) with a convex optimization problem.
@@ II-B Linear disturbance feedback policy
II-C On the choice of the clairvoyant benchmark policy
@@ III Main Results
we denote the optimal value of (III) by $\bar{\mathtt{R}}^{\star}(\bm{\Psi}_{u}(\bm{\theta}))$. Despite we narrowed attention to linear disturbance feedback policies, (III) remains intractable if $\bm{\Theta}$ has infinite cardinality. Besides, strong duality results do not apply in a straightforward way as $\bm{\Theta}$ is not assumed to be connected, let alone convex.
Motivated by the scenario optimization framework [18, 19, 20], we therefore propose replacing the maximization over $\bm{\Theta}$ with a maximization over the finite set $\mathcal{D}=\{\bm{\theta}^{1},\dots,\bm{\theta}^{N}\}$ of randomly sampled uncertainty realizations only. Proceeding in this way, we approximate (III) with its scenario counterpart, that is:
@@
with $\bm{H}_{x}^{k}=\bm{H}_{x}(\bm{\theta}^{k})$, $\bm{H}_{u}^{k}=\bm{H}_{u}(\bm{\theta}^{k})$, and $\bm{h}^{k}=\bm{h}(\bm{\theta}^{k})$ for brevity. In particular, note that the infimum in (10a) is attained since only a finite number of uncertainty realizations $\bm{\theta}^{k}\in\mathcal{D}$ are considered, and since, for every $\bm{\theta}^{k}$, $\bm{R}\succ 0$ implies that the regret (3) is radially unbounded with respect to $\bm{\Phi}_{u}$. Building upon the reformulations proposed in [9, 10] for the case of known system dynamics, the next proposition shows that (III) can be solved by means of standard convex optimization techniques.
@@
The scenario optimization problem (III) is equivalent to the following semidefinite program:
@@
$\quad\left\lVert(\bm{H}_{x}^{k}\bm{\Phi}_{x}(\bm{\theta}^{k})+\bm{H}_{u}^{k}\bm{\Phi}_{u})_{i}\bm{H}_{w}^{k}(\bm{\theta})\right\rVert_{2}\leq\bm{h}^{k}\,,$ (11b)
$\quad\begin{bmatrix}\bm{I}&\begin{bmatrix}\bm{Q}^{\frac{1}{2}}\bm{\Phi}_{x}(\bm{\theta}^{k})\\ \bm{R}^{\frac{1}{2}}\bm{\Phi}_{u}\end{bmatrix}\\ \star&\gamma\bm{I}+\begin{bmatrix}\bm{Q}^{\frac{1}{2}}\bm{\Psi}_{x}(\bm{\theta}^{k})\\ \bm{R}^{\frac{1}{2}}\bm{\Psi}_{u}(\bm{\theta}^{k})\end{bmatrix}^{\top}\begin{bmatrix}\bm{Q}^{\frac{1}{2}}\bm{\Psi}_{x}(\bm{\theta}^{k})\\ \bm{R}^{\frac{1}{2}}\bm{\Psi}_{u}(\bm{\theta}^{k})\end{bmatrix}\end{bmatrix}\succeq 0\,,$ (11k)
where $\bm{H}_{w}^{k}=\bm{H}_{w}(\bm{\theta}^{k})$, $S$ is the number of constraints in (5), and normal-$\star$ denotes entries that can be inferred from symmetry.
We remark that, for each $\bm{\theta}^{k}\in\mathcal{D}$, the operators $\bm{\Psi}_{x}(\bm{\theta}^{k})$ and $\bm{\Psi}_{u}(\bm{\theta}^{k})$ in (11k) are the noncausal system responses associated with a benchmark policy that is optimal for the specific realization $\bm{\theta}^{k}$ of the uncertain system parameters. For each $\bm{\theta}^{k}\in\mathcal{D}$, enforcing (11k) hence requires to first evaluate the corresponding optimal closed-loop behavior in hindsight using (8). Establishing regret guarantees relative to the clairvoyant optimal policy $\bm{\Psi}_{u}(\bm{\theta}^{k})$, which is impossible to compute without knowledge of $\bm{\theta}^{k}$, constitutes our main motivation towards adopting sampling techniques in a competitive setting, shedding light on an interesting application of scenario optimization beyond those in stochastic model predictive control [31, 32].
@@
As the control horizon $T$ and the number of uncertainty samples in $\mathcal{D}$ increase, solving (1) via semidefinite programming may represent a major computational bottleneck. Indeed, (1) features $\delta=1+\frac{m(T-1)(2n+p(T-2))}{2}$ optimization variables and $N$ linear matrix inequality constraints of the form (11k). As common in the predictive control literature [33, 34, 35] and inspired by time-invariant infinite horizon formulations, the computational burden can be alleviated by imposing a Toeplitz block structure on $\bm{\Phi}_{u}$, effectively reducing the number of optimization variables to $\hat{\delta}=1+m(n+p(T-2))$ – a linearly growing function of $T$. We will return to this point in Section IV, where we numerically show that this additional structure can substantially reduce the computational time, at the price of an only slight increase in conservativeness in our regret bound. We also refer the interested reader to [36, 37] for state-of-the-art techniques that leverage diagonal dominance and chordal sparsity to further improve scalability.
Let $\bm{\Phi}_{u}^{\star}(\bm{\Psi}_{u}(\bm{\theta}),\mathcal{D})$ and $\bar{\mathtt{R}}_{N}^{\star}(\bm{\Psi}_{u}(\bm{\theta}),\mathcal{D})$ denote the optimal policy and the optimal value of (III), respectively. [^7] In the interest of readability, in the following, we omit function arguments when clear from the context. Since only a finite subset of the constraints of (III) are considered in (III), we have that $\bar{\mathtt{R}}_{N}^{\star}\leq\bar{\mathtt{R}}^{\star}$, that is, $\bar{\mathtt{R}}_{N}^{\star}$ is an optimistic lower bound on the true minimax regret $\bar{\mathtt{R}}^{\star}$. Conversely, thanks to Proposition 1 and exploiting key results in scenario optimization, we now show that the solution of (III) is approximately feasible for (III) – in the sense that the measure of the set of original constraints that it violates rapidly approaches zero as $N$ increases. Before formalizing this generalization property in the theorem below, we observe that multiple optimal policies for (1) may exist, since the function $\lambda_{\max}(\cdot)$ is not strongly convex. In this case, uniqueness of $\bm{\Phi}_{u}^{\star}(\bm{\Psi}_{u}(\bm{\theta}),\mathcal{D})$ can be enforced by designing a convex tie-break rule, e.g., a lexicographic criterion [38]. Conversely, if the safety constraints (10c) are overly restrictive, the scenario problem (1) may become infeasible; if this were the case, however, the original problem (III) would also certainly be infeasible, and one would need to consider broader classes of policies, or to relax the safety requirements, e.g., by introducing slack variables in (10c).
@@
Fix any violation and confidence levels, say $\epsilon$ and $\beta$, in the open interval $(0,1)$, and let $\delta$ and $\mathbb{P}_{\bm{\theta}}^{N}$ denote the number of optimization variables in (III) and the $N$-fold product distribution $\mathbb{P}_{\bm{\theta}}\times\dots\times\mathbb{P}_{\bm{\theta}}$ with $N$ terms, respectively. If the scenario optimization problem (III) is feasible and $N>\delta$ satisfies $\sum_{j=0}^{\delta-1}\binom{N}{j}\epsilon^{j}(1-\epsilon)^{N-j}\leq\beta$, then, with probability of at least $1-\beta$ given a dataset $\mathcal{D}\sim\mathbb{P}_{\bm{\theta}}^{N}$, it holds that:
@@
Theorem 1 presents an explicit sample complexity bound that, given a priori specified $\epsilon$ and $\beta$, ensures that the safety and regret guarantees extend to all but at most a fraction $\epsilon$ of unseen dynamics $\bm{\theta}\in\bm{\Theta}$ with probability $1-\beta$. As well-known in the literature on scenario optimization, the minimum number of scenarios $N(\epsilon,\beta)$ required to fulfill the conditions of Theorem 1 grows linearly with $\epsilon^{-1}$, yet at most logarithmically with $\beta^{-1}$. Hence, even if a very small $\beta$ is selected – so that (12) holds with practical certainty – the number of scenarios to be sampled can remain manageable, see also [31]. Further, we note that the condition on the number $N$ of uncertainty samples given in Theorem 1 is tight for fully-supported problems [19]; a simpler, albeit not tight, sufficient condition on $N$ is given by [18]:
@@
III-A Comparison with worst-case oriented synthesis
@@ IV Numerical Results
In Figure 2, we compare the regret certificates $\bar{\mathtt{R}}^{\star}_{N}$ and $\hat{\mathtt{R}}^{\star}_{N}$ provided by the control policies $\bm{\Phi}_{u}^{\star}$ and $\widehat{\bm{\Phi}}_{u}^{\star}$, respectively, as well as the computation times $\bar{\tau}_{N}$ and $\hat{\tau}_{N}$ required to evaluate them via semidefinite optimization. [^9] All optimization problems have been solved using MOSEK [39] on a standard laptop computer with a 2.3 GHz Intel Core i9 CPU. Besides validating our theoretical results, these figures allow us to draw the following observations. First, the approximate solution $\widehat{\bm{\Phi}}_{u}^{\star}$ with constant block diagonal terms guarantees regret at most $9\%$ higher than $\bm{\Phi}_{u}^{\star}$ with high probability, yet its evaluation requires a computation time $\hat{\tau}_{N}$ that is lower than $\bar{\tau}_{N}$ by an entire order of magnitude. Second, consistently with the intuition that simpler models are less prone to overfit, we observe that $\widehat{\bm{\Phi}}_{u}^{\star}$ achieves better generalization than $\bm{\Phi}_{u}^{\star}$, as the out-of-sample empirical violation probability $\widehat{V}_{N}$ is consistently smaller than $\bar{V}_{N}$. Third, the quantities $\widehat{V}_{N}$ and $\hat{\mathtt{R}}^{\star}_{N}$ rapidly converge to their corresponding limit values as the number $N$ of sampled scenarios increases, suggesting that the minimax solution to (III) could be practically approximated by sampling a limited number of uncertainty instances only. Motivated by these considerations and with the aim of further reducing the computational complexity of our scheme, we plan to study the possible application of wait-and-judge [38] and constraint removal [40] approaches in future work.
Next, to illustrate the potential of our synthesis method, we compare the performance of the control policies $\bm{\pi}_{\mathtt{R}}$ and $\bm{\pi}_{\mathtt{H}}$ computed solving (1) and (14), respectively, using $N=5000$ random samples of $\delta_{k}$ and $\delta_{c}$. Then, for several classes of disturbances $\bm{w}$ often encountered in practice, we evaluate the realized control costs $J(\bm{\pi}_{\mathtt{R}},\bm{w},\bm{\theta})$, $J(\bm{\pi}_{\mathtt{H}},\bm{w},\bm{\theta})$ and $J(\bm{\psi},\bm{w},\bm{\theta})$ for $20$ different values of $\bm{\theta}$. In Figure 3a, we plot $J(\bm{\Psi}_{u}(\bm{\theta}),\bm{w},\bm{\theta})$ and compare it with $\bar{\mathtt{H}}^{\star}_{N}-\bar{\mathtt{R}}^{\star}_{N}$ to verify, according to (17), when (15) yields tighter upper bounds than (16) on the realized performance. Then, in Figure 3b, we display the percentage increase in the realized cost due to using $\bm{\pi}_{\mathtt{H}}$ instead of $\bm{\pi}_{\mathtt{R}}$, that is, [^10] For stochastic disturbances, results are averaged over $10^{4}$ realizations.

### v3

@@ Abstract
keywords:
predictive control , stochastic systems , regret minimization , scenario optimization
journal: European Journal of Control
[…]
[1]organization=Institute of Mechanical Engineering, EPFL,country=Switzerland
[…]
[2]organization=Department of Information Technology and Electrical Engineering, ETH Zürich,country=Switzerland
1 Introduction
@@
For the case of known cost functions, the formulation of [5, 6, 7] has received increasing interest thanks to: optimality of the clairvoyant benchmark policy, possibility of computing the regret-minimizing controller, and remarkable performance reported in several applications, including longitudinal motion control of a helicopter and control of a wind energy conversion system [8]. In particular, among recent contributions, [9] and [10] proposed an efficient optimization-based synthesis framework to incorporate safety constraints, [11] established recursive feasibility and stability guarantees for receding horizon regret optimal control, [12] and [13] considered partially-observed systems, [14] and [8] investigated the closely related metric of competitive ratio, [7] and [15] considered state estimation problems, [16] studied connections with imitation learning, and [17] introduced the notion of spatial regret for distributed control design.
@@
For several control applications, including robotics, building energy management, and power grids [27], designing a single state feedback policy that attains robust performance across all admissible system dynamics can prove overly conservative. Instead, it is beneficial to optimize for a unique closed-loop behavior – while allowing the state feedback law that achieves it to vary – leveraging a posteriori measurements of exogenous perturbations such as external forces, solar radiation, and electricity demands for control implementation.
@@
2 Problem Statement and Preliminaries
2.1 Dynamics, control objective, and constraints
@@
Often times, the probability distribution $\mathbb{P}_{\bm{\theta}}$ is unknown, yet uncertainty samples are directly made available to the policy designer as observations. For instance, this is the case when the realizations $\bm{\theta}^{k}\in\mathcal{D}$ correspond to a series of system identification experiments, see, e.g., [19] and [29], or to the values of the physical parameters of a batch of components, which scatter around their nominal value due to tolerance levels and variability in the production process. We refer the interested reader to [30] and [31] for examples of application of the proposed uncertainty description – which lies at the core of stochastic model predictive control [32, 33] – to wind turbine and building climate control problems, respectively.
@@
As previously discussed, measurable disturbance processes arise in several control applications and include, e.g., reference changes, friction, and external forces in robotics, and heat and humidity loads in heating, ventilating and air conditioning systems. Besides, our formulation encompasses the broad class of linear-parameter-varying systems, see [34] for a comprehensive overview and discussion of applications in automotive systems, aircraft technology, and robotics among others. In fact, if $\theta_{t}$ denotes an a priori uncertain but measurable scheduling parameter, then past disturbance realizations can always be reconstructed by $w_{t}=E_{t}(\theta_{t})^{\dagger}(x_{t+1}-A_{t}(\theta_{t})x_{t}-B_{t}(\theta_{t})u_{t})$, where $E_{t}(\theta_{t})^{\dagger}$ is the Moore-Penrose inverse of $E_{t}(\theta_{t})$. Differently from most literature on linear-parameter-varying systems, however, we allow generic nonlinear dependence with respect to $\theta_{t}$ of the system matrices $A_{t}(\theta_{t})$, $B_{t}(\theta_{t})$, and $E_{t}(\theta_{t})$.
@@
2.2 Linear disturbance feedback policy
In general, it is well-known that optimizing over the function space of feedback policies is computationally intractable. Therefore, as common in the control literature [35, 36], throughout this paper we restrict our attention to linear disturbance feedback policies of the form $\bm{u}=\bm{\Phi}_{u}\bm{w}$, with $\bm{\Phi}_{u}$ lower block-triangular to enforce causality. [^7] By carefully adapting the convex reformulation of Proposition 1 below, our results can be extended to the case of affine policies $\bm{u}=\bm{\Phi}_{u}\bm{w}+\bm{v}$ through the definition of an augmented disturbance vector $\bm{\delta}=(\bm{w},1)$. Note that linear policies attain minimum regret against the optimal sequence of control actions in hindsight if the system dynamics are known and the safety constraints are not active [5, 6, 7]. Moreover, as we will show in the next section, unlike linear state feedback policies, this choice allows us to approximate the intractable minimization of (4) subject to the dynamics (1) and the constraints (5) with a convex optimization problem.
@@ 2.2 Linear disturbance feedback policy
2.3 On the choice of the clairvoyant benchmark policy
@@ 3 Main Results
we denote the optimal value of (9) by $\bar{\mathtt{R}}^{\star}(\bm{\Psi}_{u}(\bm{\theta}))$. Despite we narrowed attention to linear disturbance feedback policies, (9) remains intractable if $\bm{\Theta}$ has infinite cardinality. Besides, strong duality results do not apply in a straightforward way as $\bm{\Theta}$ is not assumed to be connected, let alone convex.
Motivated by the scenario optimization framework [19, 20, 21], we therefore propose replacing the maximization over $\bm{\Theta}$ with a maximization over the finite set $\mathcal{D}=\{\bm{\theta}^{1},\dots,\bm{\theta}^{N}\}$ of randomly sampled uncertainty realizations only. Proceeding in this way, we approximate (9) with its scenario counterpart, that is:
@@
with $\bm{H}_{x}^{k}=\bm{H}_{x}(\bm{\theta}^{k})$, $\bm{H}_{u}^{k}=\bm{H}_{u}(\bm{\theta}^{k})$, and $\bm{h}^{k}=\bm{h}(\bm{\theta}^{k})$ for brevity. In particular, note that the infimum in (10a) is attained since only a finite number of uncertainty realizations $\bm{\theta}^{k}\in\mathcal{D}$ are considered, and since, for every $\bm{\theta}^{k}$, $\bm{R}\succ 0$ implies that the regret (3) is radially unbounded with respect to $\bm{\Phi}_{u}$. Building upon the reformulations proposed in [9, 10] for the case of known system dynamics, the next proposition shows that (10) can be solved by means of standard convex optimization techniques.
@@
The scenario optimization problem (10) is equivalent to the following semidefinite program:
@@
$\quad\left\lVert(\bm{H}_{x}^{k}\bm{\Phi}_{x}(\bm{\theta}^{k})+\bm{H}_{u}^{k}\bm{\Phi}_{u})_{i}\bm{H}_{w}^{k}\right\rVert_{2}\leq\bm{h}^{k}\,,$ (11b)
$\quad\begin{bmatrix}\bm{I}&\begin{bmatrix}\bm{Q}^{\frac{1}{2}}\bm{\Phi}_{x}(\bm{\theta}^{k})\\ \bm{R}^{\frac{1}{2}}\bm{\Phi}_{u}\end{bmatrix}\\ \star&\gamma\bm{I}+\begin{bmatrix}\bm{Q}^{\frac{1}{2}}\bm{\Psi}_{x}(\bm{\theta}^{k})\\ \bm{R}^{\frac{1}{2}}\bm{\Psi}_{u}(\bm{\theta}^{k})\end{bmatrix}^{\top}\begin{bmatrix}\bm{Q}^{\frac{1}{2}}\bm{\Psi}_{x}(\bm{\theta}^{k})\\ \bm{R}^{\frac{1}{2}}\bm{\Psi}_{u}(\bm{\theta}^{k})\end{bmatrix}\end{bmatrix}\succeq 0\,,$ (11c)
where $\bm{H}_{w}^{k}=\bm{H}_{w}(\bm{\theta}^{k})$, $S$ is the number of constraints in (5), and $\star$ denotes entries that can be inferred from symmetry.
We remark that, for each $\bm{\theta}^{k}\in\mathcal{D}$, the operators $\bm{\Psi}_{x}(\bm{\theta}^{k})$ and $\bm{\Psi}_{u}(\bm{\theta}^{k})$ in (11c) are the noncausal system responses associated with a benchmark policy that is optimal for the specific realization $\bm{\theta}^{k}$ of the uncertain system parameters. For each $\bm{\theta}^{k}\in\mathcal{D}$, enforcing (11c) hence requires to first evaluate the corresponding optimal closed-loop behavior in hindsight using (8). Establishing regret guarantees relative to the clairvoyant optimal policy $\bm{\Psi}_{u}(\bm{\theta}^{k})$, which is impossible to compute without knowledge of $\bm{\theta}^{k}$, constitutes our main motivation towards adopting sampling techniques in a competitive setting, shedding light on an interesting application of scenario optimization beyond those in stochastic model predictive control [32, 33].
@@
Differently from the computationally efficient state-space representations of a regret optimal controller available for the case where the system dynamics are known and no safety constraints are imposed on the system [5, 6, 7], our solution to the robust regret minimization problem relies on convex optimization. As the control horizon $T$ and the number of uncertainty samples in $\mathcal{D}$ increase, solving (11) via semidefinite programming may represent a major computational bottleneck. Indeed, (11) features $\delta=1+\frac{m(T-1)(2n+p(T-2))}{2}$ optimization variables and $N$ linear matrix inequality constraints of the form (11c). As common in the predictive control literature [37, 38, 39] and inspired by time-invariant infinite horizon formulations, the computational burden can be alleviated by imposing a Toeplitz block structure on $\bm{\Phi}_{u}$, effectively reducing the number of optimization variables to $\hat{\delta}=1+m(n+p(T-2))$ – a linearly growing function of $T$. We will return to this point in Section 4, where we numerically show that this additional structure can substantially reduce the computational time, at the price of an only slight increase in conservativeness in our regret bound (cf. Figure 2 in the numerical results section). We also refer the interested reader to [40, 41] for state-of-the-art techniques that leverage diagonal dominance and chordal sparsity to further improve scalability.
Let $\bm{\Phi}_{u}^{\star}(\bm{\Psi}_{u}(\bm{\theta}),\mathcal{D})$ and $\bar{\mathtt{R}}_{N}^{\star}(\bm{\Psi}_{u}(\bm{\theta}),\mathcal{D})$ denote the optimal policy and the optimal value of (10), respectively. [^8] In the interest of readability, in the following, we omit function arguments when clear from the context. Since only a finite subset of the constraints of (9) are considered in (10), we have that $\bar{\mathtt{R}}_{N}^{\star}\leq\bar{\mathtt{R}}^{\star}$, that is, $\bar{\mathtt{R}}_{N}^{\star}$ is an optimistic lower bound on the true minimax regret $\bar{\mathtt{R}}^{\star}$. Conversely, thanks to Proposition 1 and exploiting key results in scenario optimization, we now show that the solution of (10) is approximately feasible for (9) – in the sense that the measure of the set of original constraints that it violates rapidly approaches zero as $N$ increases. Before formalizing this generalization property in the theorem below, we observe that multiple optimal policies for (11) may exist, since the function $\lambda_{\max}(\cdot)$ is not strongly convex. In this case, uniqueness of $\bm{\Phi}_{u}^{\star}(\bm{\Psi}_{u}(\bm{\theta}),\mathcal{D})$ can be enforced by designing a convex tie-break rule, e.g., a lexicographic criterion [42]. Conversely, if the safety constraints (10c) are overly restrictive, the scenario problem (11) may become infeasible; if this were the case, however, the original problem (9) would also certainly be infeasible, and one would need to consider broader classes of policies, or to relax the safety requirements, e.g., by introducing slack variables in (10c).
@@
Fix any violation and confidence levels, say $\epsilon$ and $\beta$, in the open interval $(0,1)$, and let $\delta$ and $\mathbb{P}_{\bm{\theta}}^{N}$ denote the number of optimization variables in (10) and the $N$-fold product distribution $\mathbb{P}_{\bm{\theta}}\times\dots\times\mathbb{P}_{\bm{\theta}}$ with $N$ terms, respectively. If the scenario optimization problem (10) is feasible and $N>\delta$ satisfies $\sum_{j=0}^{\delta-1}\binom{N}{j}\epsilon^{j}(1-\epsilon)^{N-j}\leq\beta$, then, with probability of at least $1-\beta$ given a dataset $\mathcal{D}\sim\mathbb{P}_{\bm{\theta}}^{N}$, it holds that:
@@
Theorem 1 presents an explicit sample complexity bound that, given a priori specified $\epsilon$ and $\beta$, ensures that the safety and regret guarantees extend to all but at most a fraction $\epsilon$ of unseen dynamics $\bm{\theta}\in\bm{\Theta}$ with probability $1-\beta$. As well-known in the literature on scenario optimization, the minimum number of scenarios $N(\epsilon,\beta)$ required to fulfill the conditions of Theorem 1 grows linearly with $\epsilon^{-1}$, yet at most logarithmically with $\beta^{-1}$. Hence, even if a very small $\beta$ is selected – so that (12) holds with practical certainty – the number of scenarios to be sampled remains manageable, see also [32]. Further, we note that the condition on the number $N$ of uncertainty samples given in Theorem 1 is tight for fully-supported problems [20]; a simpler, albeit not tight, sufficient condition on $N$ is given by [19]:
@@
3.1 Comparison with worst-case oriented synthesis
@@ 4 Numerical Results
In Figure 2, we compare the regret certificates $\bar{\mathtt{R}}^{\star}_{N}$ and $\hat{\mathtt{R}}^{\star}_{N}$ provided by the control policies $\bm{\Phi}_{u}^{\star}$ and $\widehat{\bm{\Phi}}_{u}^{\star}$, respectively, as well as the computation times $\bar{\tau}_{N}$ and $\hat{\tau}_{N}$ required to evaluate them via semidefinite optimization. [^10] All optimization problems have been solved using MOSEK [43] on a standard laptop computer with a 2.3 GHz Intel Core i9 CPU. Besides validating our theoretical results, these figures allow us to draw the following observations. First, the approximate solution $\widehat{\bm{\Phi}}_{u}^{\star}$ with constant block diagonal terms guarantees regret at most $9\%$ higher than $\bm{\Phi}_{u}^{\star}$ with high probability, yet its evaluation requires a computation time $\hat{\tau}_{N}$ that is lower than $\bar{\tau}_{N}$ by an entire order of magnitude. Second, consistently with the intuition that simpler models are less prone to overfit, we observe that $\widehat{\bm{\Phi}}_{u}^{\star}$ achieves better generalization than $\bm{\Phi}_{u}^{\star}$, as the out-of-sample empirical violation probability $\widehat{V}_{N}$ is consistently smaller than $\bar{V}_{N}$. Third, the quantities $\widehat{V}_{N}$ and $\hat{\mathtt{R}}^{\star}_{N}$ rapidly converge to their corresponding limit values as the number $N$ of sampled scenarios increases, suggesting that the minimax solution to (9) could be practically approximated by sampling a limited number of uncertainty instances only. Motivated by these considerations and with the aim of further reducing the computational complexity of our scheme, we plan to study the possible application of wait-and-judge [42] and constraint removal [44] approaches in future work.
Next, to illustrate the potential of our synthesis method, we compare the performance of the control policies $\bm{\pi}_{\mathtt{R}}$ and $\bm{\pi}_{\mathtt{H}}$ computed solving (11) and (14), respectively, using $N=5000$ random samples of $\delta_{k}$ and $\delta_{c}$. For several classes of disturbances $\bm{w}$ often encountered in practice, we evaluate the realized control costs $J(\bm{\pi}_{\mathtt{R}},\bm{w},\bm{\theta})$, $J(\bm{\pi}_{\mathtt{H}},\bm{w},\bm{\theta})$ and $J(\bm{\psi},\bm{w},\bm{\theta})$ for $20$ different values of $\bm{\theta}$. In Figure 3a, we plot $J(\bm{\Psi}_{u}(\bm{\theta}),\bm{w},\bm{\theta})$ and compare it with $\bar{\mathtt{H}}^{\star}_{N}-\bar{\mathtt{R}}^{\star}_{N}$ to verify, according to (17), when (15) yields tighter upper bounds than (16) on the realized performance. In Figure 3b, we instead display the percentage increase in the realized cost due to using $\bm{\pi}_{\mathtt{H}}$ instead of $\bm{\pi}_{\mathtt{R}}$, that is, [^11] For stochastic disturbances, results are averaged over $10^{4}$ realizations.